mercvriii:

Explain how to write a linear equation in standard form using the points (1,3) and (6,-7).

izuwu1:

believe in yourself

mercvriii:

@izuwu1 ?

mercvriii:

@izuwu1 Haha, thanks. Though, I'm just stuck on this question.

Shadow:

Solve for the slope using: \[m = \frac{ y_{2} - y_{1} }{ x_{2} - x_{1} }\] Then input that into point slope form Point-Slope Form \[y - y_{1} = m(x - x_{1})\] This simplifies to: \[y = mx + b\] Which you can then convert into Standard Form: \[Ax + By = C\]

Shadow:

There might be an easier way but this way is kind of easy if you know what to do

mercvriii:

@Shadow Thanks, this helped. :)